A G's production of 950 manufactured parts contains 70 parts that do not meet the customer requirements. Two parts are selected randomly without replacement from the batch. What is the probability that the second part is defective given that the first part is defective?


1
Expert's answer
2021-05-02T08:26:23-0400

If the first part is defective, then there are "70-1=69" defective parts among "950-1=949" manufactured parts.


"P(second\ D|first\ D)=\dfrac{69}{949}"

The probability that the second part is defective given that the first part is defective is "\dfrac{69}{949}\approx0.0727."